# Intensity doubling for Brownian loop-soups in high dimensions

### v2

@@ Abstract.
We derive an intensity doubling feature of critical Brownian loop-soups on the cable-graphs of ${\mathbb{Z}}^{d}$ for $d\geq 7$ that can be described as follows: In the box $N,N]^{d}$ (and with a probability that goes to $1$ as $N$ goes to infinity), the set of all clusters of Brownian loops that do contain proper self-avoiding cycles of diameter comparable to $N$ can be decomposed into two identically distributed families: (a) The collection of clusters that do contain a large Brownian loop from the loop-soup (and therefore do automatically contain such a large cycle) (b) The collection of clusters that contain no macroscopic loop from the loop-soup (more specifically, no loop of diameter greater than $N^{\beta}$ when $\beta>4/(d-2)$ is fixed) but nevertheless contain a large cycle. In particular, due to the fact that these two families are asymptotically identically distributed, large cycles formed in case (b) by chains of small Brownian loops (i.e., all of diameter much smaller than $N$) will look like large Brownian loops themselves, and form a second independent “ghost” critical loop-soup in the scaling limit. Reformulated in terms of the Gaussian free field on such cable-graphs, this shows that large cycles in the collection of its sign clusters will converge in the scaling limit to a Brownian loop-soup with twice the “usual” critical intensity. This result had been conjectured in [16]; our proof builds heavily on the recently derived switching property for such critical cable-graph loop-soups from [21].
@@ 1. Introduction
In transient cable graphs such as those of ${\mathbb{Z}}^{d}$ for $d\geq 3$, one can define a natural Brownian loop measure and perform a Poisson point process of such Brownian loops with intensity given by this particular measure. This “Brownian loop-soup” (in the terminology of [11]) turns out to be closely related to the Gaussian Free Field (GFF) in that space (see [12, 13]) and to questions about the determinant of the Laplacian (see [13, 9, 22, 14] and the references therein). There are numerous questions that one can (try to) address about this natural random structure. The clusters of Brownian loops (the connected components of the union of all the loops in the loop-soup) are of particular importance (as for instance shown by the rewiring properties from [19] that indicate that the way in which clusters are subdivided/covered into loops is in some sense “uniform and local” or because of their relation to the Gaussian free field). Properties that will be particularly useful in the present paper are the explicit formula for the probability that two given points $x$ and $y$ belong to the same cluster of Brownian loops (we denote this event by $x\leftrightarrow y$) in terms of the Green’s function $g(x,y)$ in the graph (see [15]) or the switching property (see [21]) that allows to describe the law of the clusters when one conditions on the event $x\leftrightarrow y$. Both these features are specific to this particular percolation-of-loops model.
We choose to first present our main result in the Brownian loop-soup framework – we will briefly come back to what it says about the GFF at the end of this introduction.
@@
When the dimension of space is high (i.e., $d\geq 7$ in the case of ${\mathbb{Z}}^{d}$), the proportion of large Brownian loops vs. small Brownian loops in a given large box will be smaller and the clusters of Brownian loops (the connected components of the union of all Brownian loops) will (or are believed to) share most features of the clusters of critical Bernoulli percolation in that spatial dimension. For instance, the probability that two given points $x$ and $y$ are in the same cluster decays like a constant times $1/|x-y|^{d-2}$ when $|x-y|\to\infty$ (this follows directly from the explicit formula mentioned above in the case of loop-percolation) – the same result is known to hold for nearest-neighbour Bernoulli percolation for $d\geq 11$ and sufficiently spread-out Bernoulli percolation for $d\geq 7$, see [8] and the references therein. There has been a rather intense activity about this loop-percolation model on cable graphs – some recent references dealing with this high-dimensional regime include [20, 3, 6, 7, 4] (see more generally the recent preprint [4] for a more complete reference list).
@@
The main goal of the present paper is to derive a novel and somewhat surprising feature of this model in such high dimensions. Before describing it, let us recall some further simple observations from [20]. Consider a critical Brownian loop-soup in $\Lambda_{N}:=[-N,N]^{d}$ (viewed as a cable graph with the edges of ${\mathbb{Z}}^{d}$ in that box) for $d\geq 7$. Then:
@@
In a nutshell, this simple argument indicates that a “typical” large cluster will be composed only of Brownian loops of diameter smaller than $N^{\alpha}$ [as we shall see, it is in fact not difficult to show that this will actually be the case for all $\alpha>4/(d-2)$]. This type of ideas/results has been used and developed in subsequent work, such as [3].
We will be interested in the special exceptional macroscopic clusters that happen to contain “proper” macroscopic cycles (i.e. large self-avoiding loops that are not close to be tree-like at macroscopic scale, in a sense that we will make precise) – to avoid confusion, we will try to use the word loops for Brownian loops in a loop-soup and the word cycles for subsets of clusters (that are not necessarily Brownian loops in the loop-soup). A macroscopic Brownian loop will typically itself contain such a macroscopic “proper cycle”, so that (combined with the fact that no cluster will contain more than one macroscopic loop) we can infer that the number of such macroscopic-cycle-containing-clusters will be at least as big as the number of macroscopic Brownian loops. The question is to describe (if they exist!) the number and shape of those macroscopic-cycle-containing-clusters that do not contain any large Brownian loop from the loop-soup (so that the cycle would have to be created by a chains of smaller loops).
The answer that we will give in this paper goes as follows (a more precise formal statement will be given in the next section): For any $d\geq 7$, if $\beta\in(4/(d-2),1)$ and $\gamma\in(2/(d-4),1)$ are fixed, then, for each $\varepsilon>0$, in the limit when $N\to\infty$, one can decompose the set of clusters in $N,N]^{d}$ that do contain “proper” (i.e., not close to tree-like) macroscopic cycles with diameter greater $\varepsilon N$ into two (almost) identically distributed ones (as depicted in Figure 1):
@@
    The ones that contain exactly one macroscopic Brownian loop of the loop-soup (and this Brownian loop will already contain a cycle) as discussed above,
@@
    the ones that do contain no Brownian loop of diameter greater than $N^{\beta}$,
and there will be no other ones (with a probability that goes to $1$ as $N\to\infty$). In fact, for Case (1), the clusters will contain no Brownian loop of diameter greater than $N^{\beta}$ other than the one of diameter larger than $\varepsilon N$ that it already contains by definition, while for Case (2), the large cycles will in fact be created by Brownian loops of diameter smaller than $N^{\gamma}$ (i.e., loops of diameter larger than $N^{\gamma}$ can be in the cluster, but will not contribute to the creation of the large cycle). One main point in this statement is that the set of clusters satisfying (1) and (2) are (almost) identically distributed. In fact, for each cluster, the choice of being of the first or second kind is made by (almost) fair independent coins (one for each cluster – we will explain that the independence follows readily from the Poissonian feature of the loop-soup); the exact way to formulate this $N\to\infty$ statement is that, with a probability that tends to $1$ as $N\to\infty$, when one conditions on all the macroscopic clusters, the conditional probabilities for each cluster to be of type (1) and (2) respectively will both be $1/2+o(1)$, and the conditional probability of being in neither case will be $o(1)$.
Figure 1. Each cluster containing a large cycle will be decomposed in one of these two ways with (almost) equal probability: Either the cycle is due to one big Brownian loop, or the cycle is created by a chain of very small Brownian loops.
@@
As explained in [15], clusters of critical Brownian loop-soups can be equivalently described as the excursions sets away from 0 of a Gaussian Free Field (GFF) on the cable graph. This makes it possible to formulate a consequence of our main result in terms of the GFF as follows: Large proper cycles in the excursions sets of the GFF in the cable graph of $N,N]^{d}$ will converge in the scaling limit to a Brownian loop-soup with intensity equal to twice that of the usual critical loop-soup in $1,1]^{d}$. This is exactly the “intensity doubling conjecture” formulated in [16] (based on some formulas for “twisted” loop-soups and Gaussian free fields).
@@ 2. The setup and the more precise statement of the main result
To state the result precisely, we need to define our set of “proper non-tree-like” cycles. One way to proceed is to fix some small $\varepsilon>0$ and to define ${\mathcal{C}}(\varepsilon,N)$ to be the set of loop-soup clusters (for the cable-graph loop-soup in $\Lambda_{N}:=[-N,N]^{d}$ for $d\geq 7$) that do contain a simple cycle $\Gamma$ such that there exists a plane $\Pi$ such that the orthogonal projection of $\Gamma$ on $\Pi$ has a non-trivial index around some point $x$ in this plane and also stays at distance at least $\varepsilon N$ from it (so in particular, this cycle $\Gamma$ has necessarily a diameter of at least $2\varepsilon N$). This definition ensures that the cycle is macroscopic and also not tree-like at macroscopic level (as it circumnavigates around an entire “tube”).
@@
The closer $\beta$ and $\gamma$ are to the values 4/(d2)4/(d-2) and 2/(d4)2/(d-4) (respectively), the stronger the statement is. Note also of course that 4/(d2)>2/(d4)4/(d-2)>2/(d-4) so that one can choose $\gamma$ smaller than 4/(d2)4/(d-2). One can indeed find $\beta$ and $\gamma$ smaller than $1$ that satisfy the conditions in the theorem as soon as $d\geq 7$, while (and this is not a surprise here) when one formally plugs in $d=6$ into the formula, the bounds 4/(d2)4/(d-2) and 2/(d4)2/(d-4) both take the critical value $1$. The range of admissible values for $\beta$ and $\gamma$ increases with $d\geq 7$. In particular, any values $\beta>4/5$ and $\gamma>2/3$ will work for all $d\geq 7$. As we will explain at the end of the paper, the values of these two thresholds 4/(d2)4/(d-2) and 2/(d2)2/(d-2) are not surprising and can be heuristically explained rather simply.

### v3

@@ Abstract.
We derive an intensity doubling feature of critical Brownian loop-soups on the cable-graphs of ${\mathbb{Z}}^{d}$ for $d\geq 7$ that can be described as follows: In the box $N,N]^{d}$ (and with a probability that goes to $1$ as $N$ goes to infinity), the set of all clusters of Brownian loops (in this Poissonian collection of such loops) that do contain proper self-avoiding cycles of diameter comparable to $N$ can be decomposed into two identically distributed families: (a) The collection of clusters that do contain a large Brownian loop from the loop-soup (and therefore do automatically contain such a large cycle) (b) The collection of clusters that contain no macroscopic loop from the loop-soup (more specifically, no loop of diameter greater than $N^{\beta}$ when $\beta\in(4/(d-2),1)$ is fixed) but nevertheless contain a large cycle. In particular, due to the fact that these two families are asymptotically identically distributed, large cycles formed in case (b) by chains of small Brownian loops (i.e., all of diameter smaller than $N^{\beta}$) will look like large Brownian loops themselves, and form a second independent “ghost” critical loop-soup in the scaling limit. Reformulated in terms of the Gaussian free field on such cable-graphs, this shows that large cycles in the collection of its sign clusters will converge in the scaling limit to a Brownian loop-soup with twice the “usual critical intensity”. This result had been conjectured in [27]; our proof builds heavily on the recently derived switching property for such critical cable-graph loop-soups from [36].
@@ 1. Introduction
Background on cable-graph loop-soups
In transient cable graphs such as those of ${\mathbb{Z}}^{d}$ for $d\geq 3$, one can define a natural Brownian loop measure and perform a Poisson point process of such Brownian loops with intensity given by this particular measure. This “Brownian loop-soup” (in the terminology of [21]) turns out to be closely related to the Gaussian Free Field (GFF) in that space (see [22, 23]) and to questions about the determinant of the Laplacian (see [23, 19, 37, 24] and the references therein).
We will briefly recall the exact definition of the Brownian loop-soup in Section 2, but let us already give some heuristic intuition about its discrete analog. If one considers the discrete graph ${\mathbb{Z}}^{d}$, it is natural to consider the measure $\nu$ that puts a mass $(2d)^{-n(\gamma)}$ to each nearest-neighbour loop $\gamma$ with $n(\gamma)$ steps (the loop is here considered to be “not rooted” i.e., with no marked starting point, and also “not oriented” so that a loop and its “time-reversal” are considered to be the same – so it a nearest-neighbour from ${\mathbb{Z}}/n{\mathbb{Z}}$ into ${\mathbb{Z}}^{d}$ defined modulo circular time-shift and time-reversal). The Brownian loop-measure $\nu_{B}$ that we will use here is the simply the Brownian analog of this measure on discrete unrooted unoriented loops, where one now considers Brownian loops on the cable graph (which is the union of all the one-dimensional edges joining neighboring points in ${\mathbb{Z}}^{d}$) instead of discrete random walk loops.
The Brownian loop-soup is then a Poisson point process of unrooted and unoriented Brownian loops with this particular intensity. In other words, random circular Brownian structures appear “independently” on the cable-graph, with intensity given by the Brownian loop-measure. This can appear somewhat convoluted at first sight, but from a “continuum” metric cable-graph perspective (where points/sites play no particular role), it is a particularly intrinsic percolation model. Indeed, Brownian paths are canonically defined from the metric and the Brownian loop measures furthermore remains the same if one multiplies the metric by any constant (which contrasts with usual Bernoulli bond-percolation on discrete graphs say, that can be viewed as coming from a Poisson point process of points/cuts on the cable graph, which is arguably the other very canonical model – in this case, multiplying the metric by a constant amounts to changing the percolation parameter).
It turns out that if one samples such a loop-soup on the cable-graph of ${\mathbb{Z}}^{d}$ for $d\geq 3$, for every given point $x$, there will almost surely exist infinitely many small loops that go through $x$ (but only finitely many loops of diameter greater than $\epsilon$, for any fixed $\epsilon$). On the other hand, it is not difficult to see there will almost surely exist many exceptional points on the cable graph that are visited by no loop of the loop-soup.
It is interesting to look also at Brownian loop-soups with intensities given by a multiple $c\not=1$ of $\nu_{B}$, but the case $c=1$ that we are focusing on in the present paper has some important special features. It is worth stressing here to avoid confusion that in the literature (for instance in [22, 25]), one often starts with the natural measure on oriented loops, which corresponds to $\tilde{\nu}_{B}=2\nu_{B}$, so that the $c=1$ case that we are considering is the one with intensity $\alpha\tilde{\nu}_{B}$ for $\alpha=1/2$.
@@
There are several related reasons that make it particularly natural and particularly physically relevant to then look at the clusters of Brownian loops (the connected components of the union of all the loops in the loop-soup) for $c=1$: One is the rewiring properties from [34] that indicate that the way in which clusters are subdivided/covered into individual loops is in some sense “uniform and local” – this can be viewed as a “bosonic feature” of the model. Another feature is the relation of the loop-soup with the Gaussian free field (GFF) and its correlations. There has been a rather intense activity on this particular model on cable graphs in this past decade – references include [1, 6, 7, 5, 9, 8, 11, 12, 13, 14, 15, 16, 17, 25, 26, 27, 28, 29, 30, 31, 34, 35, 36]. Results that will be particularly useful in the present paper are the explicit formula for the probability that two given points $x$ and $y$ belong to the same cluster of Brownian loops (we denote this event by $x\leftrightarrow y$) in terms of the Green’s function $g(x,y)$ in the graph (see [25]) or the switching property (see [36]) that allows to describe the law of the clusters when one conditions on the event $x\leftrightarrow y$. Both these features are specific to this particular percolation-of-loops model. One can interpret these specific features as part of the reason why this percolation model is physically natural, or as the technical reason for which one ends up being able to say more about it than for ordinary percolation. We choose to first present our main result in the Brownian loop-soup framework – we will briefly come back to what it says about the GFF at the end of this introduction.
@@
These loop-soup clusters are of interest in any dimension (in the planar case, this ends up being part of the conformal invariance SLE/CLE realm, see for instance [32]). The focus of the present paper with be the case where the dimension of space is high (i.e., $d\geq 7$ in the case of ${\mathbb{Z}}^{d}$). There, the proportion of large Brownian loops versus small Brownian loops in a given large box will be smaller and the clusters of Brownian loops (the connected components of the union of all Brownian loops) will (or are believed to) share most features of the clusters of critical Bernoulli percolation in that spatial dimension. For instance, the probability that two given points $x$ and $y$ are in the same cluster decays like a constant times $1/|x-y|^{d-2}$ when $|x-y|\to\infty$ (this follows directly from the explicit formula mentioned above in the case of loop-percolation) – the same result is known to hold for nearest-neighbour Bernoulli percolation for $d\geq 11$ and sufficiently spread-out Bernoulli percolation for $d\geq 7$, see [18] and the references therein. Some recent references dealing with this high-dimensional regime of Brownian loop-percolation include [35, 5, 15, 16, 9] (see more generally the recent paper [9] for a more complete reference list).
Loop-soups in high dimensions
@@
The main goal of the present paper is to derive a novel and somewhat surprising feature of this model in such high dimensions that also provides new insight into critical models in high-dimensions in general. Before describing it, let us recall some further simple observations from [35]. Consider a critical Brownian loop-soup in $\Lambda_{N}:=[-N,N]^{d}$ (viewed as a cable graph with the edges of ${\mathbb{Z}}^{d}$ in that box) for $d\geq 7$. More precisely, this is the collection all the Brownian loops in a loop-soup in ${\mathbb{Z}}^{d}$ that stay in $\Lambda_{N}$ (it can also be defined directly in $\Lambda_{N}$ where the boundary of the box plays the role of killing points for the Brownian motion). Then:
@@
In a nutshell, this simple argument indicates that a “typical” large cluster will be composed only of Brownian loops of diameter smaller than $N^{\alpha}$ [as we shall see, it is in fact not difficult to show that this will actually be the case for all $\alpha>4/(d-2)$]. This type of ideas/results has been used and developed in subsequent recent work, such as [5].
Results of this paper
We will be interested in the special exceptional macroscopic clusters that happen to be topologically non-trivial at macroscopic scale, i.e., that contain “proper” macroscopic cycles (i.e. large self-avoiding cycles that are not close to be tree-like at macroscopic scale, in a sense that we will make precise) – to avoid confusion, we will try to use the word loops for Brownian loops in a loop-soup and the word cycles for subsets of clusters (that are not necessarily Brownian loops in the loop-soup). A macroscopic Brownian loop will typically itself contain such a macroscopic “proper cycle”, so that (combined with the fact that no cluster will contain more than one macroscopic loop) we can infer that the number macroscopically topologically non-trivial clusters will be at least as big as the number of macroscopic Brownian loops. The question is to describe (if they exist!) the number and shape of those large cycle-containing-clusters that do not contain any large Brownian loop from the loop-soup (so that the cycle would have to be created by a chains of smaller loops).
In a nutshell, the answer will be that such clusters will exist, and that they will be as numerous and have a similar shape as the set of clusters that do contain a large Brownian loop. In somewhat less loose terms (a precise formal statement will be given in the next section): For any $d\geq 7$, if $\beta\in(4/(d-2),1)$ is fixed, then, for each $\varepsilon>0$, in the limit when $N\to\infty$, one can decompose the set of clusters in $N,N]^{d}$ that do contain “proper” (i.e., not close to tree-like) macroscopic cycles with diameter greater $\varepsilon N$ into two (almost) identically distributed families:
@@
    The set of clusters that contain exactly one macroscopic Brownian loop of the loop-soup (and this Brownian loop will already contain a cycle, which will be essentially the only cycle present in the cluster) as discussed above – this case is depicted on the left of Figure 1,
@@
    the set of clusters that do contain no Brownian loop of diameter greater than $N^{\beta}$ (as depicted on the right of Figure 1),
and there will be no other ones (with a probability that goes to $1$ as $N\to\infty$).
One main point in this statement is that the set of clusters satisfying (1) and (2) are (almost) identically distributed. So, the number of clusters of type (2) will converge to a Poisson random variable and chains of loops of size smaller than $N^{\beta}$ will form a collection of cycles that will look like a Poisson cloud of macroscopic Brownian loops.
In fact, for each individual cluster, the choice of being of the first or second kind is made by (almost) fair independent coins (one for each cluster – we will explain that the independence follows readily from the Poissonian feature of the loop-soup); a more precise way to formulate this $N\to\infty$ statement is that, with a probability that tends to $1$ as $N\to\infty$, when one conditions on all these macroscopic clusters, the conditional probabilities for each cluster that contains a macroscopic cycle to be of type (1) and (2) respectively will both be $1/2+o(1)$, and the conditional probability of being in neither case will be $o(1)$.
In fact, for Case (1), the clusters will also contain no Brownian loop of diameter greater than $N^{\beta}$ other than the one of diameter larger than $\varepsilon N$ that it already contains by definition of this case. In Case (2), the large cycles will in fact be created by Brownian loops of diameter much smaller than $N^{\beta}$ (i.e., with probability going to $1$, they will be created by loops smaller than $N^{\gamma}$ for any given $\gamma\in(2/(d-4),\beta)$ (i.e., loops of diameter larger than $N^{\gamma}$ but smaller than $N^{\beta}$ can be in the cluster, but will not contribute to the creation of the large cycle).
Figure 1. Each cluster topologically non-trivial cluster containing a large cycle will be decomposed in one of these two ways with (almost) equal probability: Either the cycle is due to one big Brownian loop (left picture), or the cycle is created by a chain of very small Brownian loops (right picture).
@@
As explained in [25], clusters of critical Brownian loop-soups can be equivalently described as the excursions sets away from 0 of a Gaussian Free Field (GFF) on the cable graph. This makes it possible to formulate a consequence of our main result in terms of the GFF as follows: Large proper cycles in the excursions sets of the GFF in the cable graph of $N,N]^{d}$ will converge in the scaling limit to a Brownian loop-soup with intensity equal to twice that of the usual critical loop-soup in $1,1]^{d}$. This is the “intensity doubling conjecture” formulated in [27] (based on some formulas for “twisted” loop-soups and Gaussian free fields) – that is therefore proved in the present paper for all $d\geq 7$.
@@
Let us finally comment about what this result says for critical models in high-dimensions in general: For many critical models from statistical physics, there is usually an “upper-critical dimension” of space above which the model will behave in a “mean-field way”. This is often related to the inter-related facts that (1) the correlation functions behave like that of the GFF, (2) the geometric underlying macroscopic clusters tend to proliferate and there is no geometric scaling limit, (3) these clusters will typically be tree-like on large scale (with no macroscopic loops). What the present paper shows is that in fact, if one looks at the exceptional clusters that are topologically non-trivial (instead of looking at all/typical clusters), scaling limits will in fact arise in the form of Brownian loop-soups. See the paper [10] (and further forthcoming work) for the case of ordinary percolation.
@@ 2. The setup and the more precise statement of the main result
As promised, we now briefly the definition of the Brownian loop-soup on transient cable-graphs: The cable-graph consists of the union of all the closed edges of the graph (viewed as one-dimensional segments), and it is then easy to define Brownian motion on this locally one-dimensional structure (when the Brownian motion is at a vertex where several edges meet, it essentially tosses a uniform coin to decide in which of the edges to move when it makes an excursion away from the vertex). From this, one can then define for each point $x$ in the cable-graph, a Brownian excursion measure $\mu_{x}$ away from $x$ in the cable graph, which is a measure on parameterized oriented loops $\gamma$ from $x$ to $x$ (we call this a loop that is rooted at xx)– a loop $\gamma$ has a positive and finite time-duration $T(\gamma)$. The total mass of $\mu_{x}$ is then the Green’s function $G(x,x)$ in the cable graph. In the set-up that we are looking at in the present paper, it is natural to consider the loops to be “non-oriented” meaning that a loop and the loop traced backwards will be considered to be the same. This makes it natural to consider the measure $\mu_{x}/2$ instead of $\mu_{x}$. The natural loop-measure is then defined by
$\mu(A):=\int_{\mathcal{G}}dx\int\frac{d\mu^{x}(\gamma)}{2}\frac{1_{\gamma\in A}}{T(\gamma)},.$
where we view $\mu$ as a measure on the set of unrooted and unoriented loops – i.e., two rooted oriented loops are identified if they are obtained by monotone circular time-reparameterization of each other) and $dx$ denotes the one-dimensional Lebesgue measure on $\mathcal{G}$}. The term $1/2$ can be seen as coming from the oriented/unoriented feature mentioned in the previous section, while the $1/T(\gamma)$ term corresponds to the idea that an unrooted loop of time-length $T$ has $T$ times more options for choosing its root than a loop of time-length $1$ (so that this term is here to compensate the natural overcounting due to that – one can view the point $x$ in the previous construction as being chosen uniformly (with respect to time) on the unrooted loop). One standard feature (going back to the the original construction from [21] in the continuum) is that it is easy to choose to “root” a loop at some special point (for instance, the point where the first coordinate is the largest) and to then describe the law of the loop viewed from that point. One can also (and we will use this in the present paper) choose some point $x$ on the cable-graph and describe the measure $\mu$ restricted to the set of loops that go through $x$ in terms of Brownian loops “rooted at xx”.
So, in a nutshell, $\mu$ is a natural measure on unoriented unrooted loops in the cable-graph, and the loop-soup that we will consider in the present paper is a Poisson point process with intensity $\mu$. Heuristically speaking, this means that loops appear independently in a loop-soup, with an “intensity” provided by $\mu$. As mentioned in the introduction, when the cable-graph is transient, it will turn out that (a) for every given point in the cable-graph, there will almost surely be infinitely many small Brownian loops in the loop-soup that go through this point, but only finitely many ones of diameter greater than $1/2$ say, and (b) there will be exceptional points that are visited by none of the loops in the loop-soup (let us call $Z$ this set). We can then define clusters of Brownian loops to be the connected components of the unions of all loops in the loop-soup (or equivalently the connected components of the complement of $Z$).
Throughout the paper, when we talk about the loop-soup in $\Lambda_{N}=[-N,N]^{d}$, we will mean a loop-soup in the cable-graph of ${\mathbb{Z}}^{d}$ restricted to the set of loops that stay in $\Lambda_{N}$ (as noted in the final section, our results can then be directly used to obtain the similar results if one considers a loop-soup in the whole of ${\mathbb{Z}}^{d}$ and considers the large cycles or the large cycles it creates that are contained in $\Lambda_{N}$.
To state the result precisely, we need to define our set of “proper non-tree-like” cycles, which is what we now do: One way to proceed is to fix some small $\varepsilon>0$ and to define ${\mathcal{C}}(\varepsilon,N)$ to be the set of loop-soup clusters (for the cable-graph loop-soup in $\Lambda_{N}:=[-N,N]^{d}$ for $d\geq 7$) that do contain a simple cycle $\Gamma$ such that there exists a plane $\Pi$ such that the orthogonal projection of $\Gamma$ on $\Pi$ has a non-trivial index around some point $x$ in this plane and also stays at distance at least $\varepsilon N$ from it (so in particular, this cycle $\Gamma$ has necessarily a diameter of at least $2\varepsilon N$). This definition ensures that the cycle is macroscopic and also not tree-like at macroscopic level (as it circumnavigates around an entire “tube”).
@@
The closer $\beta$ and $\gamma$ are to the values 4/(d2)4/(d-2) and 2/(d4)2/(d-4) (respectively), the stronger the statements (1)-(2)-(3) are. The statement is not void since one can always find $\beta$ and $\gamma$ smaller than $1$ that satisfy the conditions in the theorem when $d\geq 7$, while (and this is not a surprise here) when one formally plugs in $d=6$ into the formula, the bounds 4/(d2)4/(d-2) and 2/(d4)2/(d-4) both take the critical value $1$. The range of admissible values for $\beta$ and $\gamma$ increases with $d\geq 7$. In particular, any values $\beta>4/5$ and $\gamma>2/3$ will work for all $d\geq 7$. As we will explain at the end of the paper, the values of these two thresholds 4/(d2)4/(d-2) and 2/(d4)2/(d-4) are not surprising and can be heuristically explained rather simply.